### v1

# Two Micron-Size Dark Dimensions

Abstract
Two extra dimensions of micron scale might simultaneously address the gauge and cosmological hierarchy problems. In our paper we examine various observational bounds in scenarios with one and two large extra dimensions, to see if they are compatible with the micron scale. We show that consistency with astrophysical and cosmological observations requires that two extra dimensions of micron scale must not admit isometries, whereby conservation of the extra dimensional momentum is violated to a large extent, allowing thus the massive Kaluza-Klein modes of the graviton to promptly decay to other lighter graviton modes. We also show that within this scenario primordial black holes with masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$ could make all cosmological dark matter.
preprint: MPP-2025-5preprint: LMU-ASC 02/25
I General Idea
Extra dimensions are in general thought to be compact, since a four-dimensional (4D) description fits the world we perceive and measure (so far) very well. It is this compactness that sets a new length scale. A particularly attractive framework has as its main feature large extra dimensions along which only gravity propagates, because Standard Model (SM) fields are localized on a brane extended in the non-compact dimensions Arkani-Hamed:1998jmv . Within this framework the usual 4D graviton is complemented by a tower of Kaluza-Klein (KK) modes, corresponding to the new available phase space in the bulk. In the canonical example, spacetime is a direct product of a non-compact 4D spacetime manifold and a flat spatial $n$-torus of common linear size $2\pi R_{\rm KK}$ and volume $V_{n}=(2\pi R_{\rm KK})^{n}$. The $(4+n)$ dimensional Planck scale (or more generally the species scale $\Lambda_{\rm sp}$ where gravity becomes strong Dvali:2007hz ; Dvali:2007wp ; Dvali:2009ks ; Dvali:2010vm ; Dvali:2012uq ) is related to the 4D reduced Planck mass $M_{p}$ according to
$\Lambda_{\rm sp}=M_{p}^{2/(2+n)}\ V_{n}^{1/(2+n)}\,.$ (1)
Most notably, this framework can be embedded into string theory Antoniadis:1998ig .
Besides, by combining this idea with conjectures of the Swampland program Vafa:2005ui ; Palti:2019pca ; vanBeest:2021lhn ; Agmon:2022thq a deep interplay between physics in the UV and the IR becomes exposed. For example, the distance conjecture states that infinite distance limits $\Delta\phi\to\infty$ in the moduli space of massless scalar fields are accompanied by an infinite tower of exponentially light states $m_{\rm KK}\sim e^{-\alpha\Delta\phi}$, where distance and masses are measured in Planck units Ooguri:2006in . Connected to the distance conjecture is the anti-de Sitter (AdS) distance conjecture, which correlates the dark energy density to the mass scale $m_{\rm KK}$ characterizing the infinite tower of states,
$m_{\rm KK}\sim|\Lambda|^{\alpha}\,,$ (2)
as the negative AdS vacuum energy $\Lambda\to 0$, where $\alpha$ is an ${\cal O}(1)$ positive constant Lust:2019zwm . Besides, under the hypothesis that this scaling behaviour (2) holds in de Sitter (or quasi de Sitter) space, an unbounded number of massless modes should also emerge in the limit $\Lambda\to 0$.
Note that the KK (or string) tower contains massive spin-2 bosons, and thus for positive $\Lambda\to 0$ the Higuchi bound Higuchi:1986py places an absolute upper limit to the exponent: $\alpha=1/2$. In addition, explicit string calculations of the vacuum energy set a lower bound: $\alpha=1/4$ Montero:2022prj (see e.g. also Itoyama:1986ei ; Itoyama:1987rc ; Antoniadis:1991kh ; Bonnefoy:2018tcp ; Burgess:2023pnk ; Branchina:2023ogv ; Anchordoqui:2023laz ; Branchina:2024ljd ; Aoufia:2024awo ; Basile:2024lcz ). Now, the combination of the theoretically accessible parameter range $21/4\leq\alpha\leq 1/$ 2 with the experimental constraints on deviations from Newton’s gravitational inverse-square law Lee:2020zjt , namely
$R_{\rm KK}\lesssim 30~{}\mu{\rm m}\,,$ (3)
selects $\alpha=1/4$ as the preferred value and gives in this way a procedure to connect $R_{\rm KK}$ to the dark energy scale $\Lambda/M_{p}^{4}\sim 10^{-120}$ (i.e. $\Lambda^{-1/4}\sim 2.39~{}{\rm meV}$ Planck:2018vyg ) as follows Montero:2022prj
$R_{\rm KK}\sim\lambda\ \Lambda^{-1/4}\simeq{\cal O}(\mu{\rm m})\,,$ (4)
where the proportionality factor is estimated to be within the range $10^{-1}<\lambda<10^{-4}$.
Actually, the relation (4) is independent of the number $n$ of dark dimensions of micron size. Taking into account the supernovae and neutron star constraints (see the discussion in the next section), the authors of Montero:2022prj have argued in favour of a single dark dimension in the micron range. In this case, one is led to a species scale of order Montero:2022prj
$\Lambda_{\rm sp}\sim M_{p}^{2/3}\ R_{\rm KK}^{-1/3}\sim M_{p}^{2/3}\ \Lambda^{1/12}\sim 10^{9}~{}{\rm GeV}\,.$ (5)
In this paper we explore the possibility of a compact space with two dark dimensions of micron scale. In this case, the species scale
$\Lambda_{\rm sp}\sim(M_{p}/R_{\rm KK})^{1/2}\sim M_{p}^{1/2}\Lambda^{1/8}$ (6)
would be in the 10 TeV regime, within reach of the Future Circular Collider in the modality of hadron-hadron interactions (FCC-hh) FCC:2018vvp . Moreover, this scenario could simultaneously account for the cosmological and gauge hierarchy problems. The goal of this article is to establish the required conditions for (6) to be consistent with experiment.
The layout of the paper is as follows. In Sec. II we confront (6) to astrophysical data and in Sec. III to cosmological observations. In Sec. IV we reexamine whether neutrino masses and mixings could exclusively occur due to physics in the bulk. In Sec. V we show that an all-dark-matter interpretation in terms of 6D primordial black holes (PBH) should be feasible for masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$. In Sec. VI we take a look at the possibility of probing $\Lambda_{\rm sp}$ with ultra-high-energy cosmic rays. Conclusions are given in Sec. VII.
II Astrophysical Constraints
Supernova (SN) explosions have long been considered to be powerful probes of large extra dimensions Arkani-Hamed:1998sfv . This is because SN cores could emit sizable fluxes of KK gravitons, with masses up to about 100 MeV. The KK emission process could then compete with neutrino cooling, shortening the observable signal. This implies that the size of the extra dimensions could be constrained by requiring that SN 1987A did not emit more KK gravitons than is compatible with the observed neutrino signal duration Cullen:1999hc ; Barger:1999jf ; Hanhart:2000er ; Hanhart:2001fx . This reasoning leads to a bound
$R_{\rm KK}<\left\{\begin{array}[]{ll}490~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 0.96~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (7)
corresponding to $\Lambda_{\rm sp}>740~{}{\rm TeV}$ and $\Lambda_{\rm sp}>8.9~{}{\rm TeV}$, respectively Hannestad:2003yd .
A more restrictive constraint emerges by considering the radiative decay of KK gravitons, which could produce potentially observable gamma rays Hannestad:2001jv . The partial decay width of this process is estimated to be
$\Gamma^{l}_{\gamma\gamma}=\frac{{\tilde{\lambda}}^{2}\ m^{3}_{\rm KK}\ l^{3}}{80\pi M_{p}^{2}}\,,$ (8)
where we have considered a tower of equally spaced gravitons, indexed by an integer $l$ and with mass $m_{l}=l\ m_{\rm KK}$, and where the parameter $\tilde{\lambda}$ measures the value of the dark graviton wave function at the SM brane and is expected to be ${\cal O}(1)$ Han:1998sg ; Hall:1999mk . The non-observation of these gamma rays in data collected by the Energetic Gamma Ray Experiment Telescope (EGRET) leads to the following limits on the compactification scale:
  • The diffuse $\gamma$-ray flux measured by the EGRET instrument constrains the number of KK-gravitons that may have been emitted by all cosmic SNe. This constraints can be translated into the bound
    $R_{\rm KK}<\left\{\begin{array}[]{ll}4.9~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 96~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (9)
    corresponding to limits of $\Lambda_{\rm sp}>3.4\times 10^{3}~{}{\rm TeV}$ and $\Lambda_{\rm sp}>28~{}{\rm TeV}$, respectively Hannestad:2003yd .
  • Cas A is a young SN remnant. Based on its age, a cloud of emitted KK-gravitons should still appear as a point source to EGRET. However, EGRET does not observe a photon flux at the expected point in space. The difference between the theoretically expected flux and observation restricts the emissivity of KK gravitons. This restriction can be translated into the bound
    $R_{\rm KK}<\left\{\begin{array}[]{ll}410~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 0.36~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (10)
    corresponding to limits of $\Lambda_{\rm sp}>770~{}{\rm TeV}$ and $\Lambda_{\rm sp}>14~{}{\rm TeV}$, respectively Hannestad:2003yd .
  • In a core collapse SN most KK gravitons are produced close to the kinematic threshold. For SN core temperature $\sim 30~{}{\rm MeV}$, the threshold condition implies that the typical mass of KK gravitons is ${\cal O}(100~{}{\rm MeV})$. Thus, most gravitons leave the SN core with rather non-relativistic velocities such that a large fraction of them ends up being gravitationally retained in a cloud around the neutron star remnant. Trapped in this cloud the KK gravitons would subsequently decay to SM particles on a time scale comparable to the age of the universe. We would then expect that neutron stars should shine brightly in 100 MeV gamma-rays. However, this is not the case relative to EGRET data, yielding
    $R_{\rm KK}<\left\{\begin{array}[]{ll}7.4~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 51~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (11)
    corresponding to limits of $\Lambda_{\rm sp}>2.9\times 10^{3}~{}{\rm TeV}$ and $\Lambda_{\rm sp}>39~{}{\rm TeV}$, respectively Hannestad:2003yd .
The decay of KK gravitons also leads to an excess heating of neutron stars Hannestad:2001xi . This heating effect should have been seen by the Hubble Space Telescope which is able to observe the thermal emission from the surface of neutron stars. The lack of such an excess heat leads to the most stringent bounds on the size of the compact space,
$R_{\rm KK}<\left\{\begin{array}[]{ll}44~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 0.15~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (12)
corresponding to $\Lambda_{\rm sp}>1.6\times 10^{5}~{}{\rm TeV}$ and $\Lambda_{\rm sp}>701~{}{\rm TeV}$, respectively Hannestad:2003yd .
Notwithstanding, the constraints from EGRET data and neutron-star heating rely on the assumption that KK modes can only decay into SM fields localized on the brane, but cannot decay into other KK modes with smaller bulk momenta. If large extra dimensions do not admit isometries this assumption is in general not valid and the bounds in (9), (10), (11), and (12) can be avoided Mohapatra:2003ah . To demonstrate this point we assume that the KK modes acquire a non-zero vacuum expectation value, $\langle\varphi_{\vec{n}}\rangle$, which is independent of $l$ and leads to a spontaneous breakdown of the translational invariance in the compact space. If this were the case, the violation of KK momentum conservation would allow intra-tower decays. The total decay width of $l$ KK graviton is found to be Mohapatra:2003ah
$\Gamma_{\rm tot}^{l}\sim\Gamma^{l}_{\gamma\gamma}\left(\frac{\langle\varphi_{\vec{n}}\rangle\ M_{p}}{\Lambda_{\rm sp}^{2}}\right)^{2}\,.$ (13)
Note that for $\langle\varphi_{\vec{n}}\rangle\sim\Lambda_{\rm sp}\sim 10~{}{\rm TeV}$, the total decay rate (13) is about 28 orders of magnitude larger than (8). As a consequence, the $l$ KK graviton would predominantly decay into lighter KK gravitons.
It was recently noted that the amount of energy that can be absorbed by the dark matter halo constrains the fraction of the supernova energy which could released via the form of dark radiation to roughly $10^{-5}$ Vogl:2024ack . However, we note that the efficiency of absorbing the free-streaming gravitons is extraordinarily small.
We conclude that in the absence of isometries the SN constraints are given by (7) and two micron size dimensions with $R_{\rm KK}\sim 1~{}\mu{\rm m}$ are possible. On the other hand, if the bounds (11) and (12) are applied, the only possible case is $n=1$ corresponding to one micron size dark dimension.
III Cosmological Constraints
Most of the cosmological constraints on large extra dimensions have also been derived on the assumption that the KK momentum of the graviton is conserved (see e.g. Hall:1999mk ; Fairbairn:2001ct ; Hannestad:2001nq ), and therefore these bounds can be avoided if the compact space is not translational invariant. We have seen that if the KK momentum of the graviton tower is not conserved a given KK mode of the tower could decay into final states that include other, lighter KK excitations. Such a “dark-to-dark” decay process usually dominates the cosmological evolution of the graviton tower and could be regarded as the telltale signature of the dynamical dark matter framework Dienes:2011ja .
A specific realization of dynamical dark matter has been proposed for $n=1$, assuming that a parent particle with mass $m_{l}$ can decay to two daughter particles with masses $m_{l^{\prime}}$ and $m_{l^{\prime\prime}}$ such that $m_{l}=m_{l^{\prime}}+m_{l^{\prime\prime}}+\epsilon$, with $\epsilon\leq m_{\rm KK}\delta$ and $\delta\sim{\cal O}(1)$ Gonzalo:2022jac . As might be expected, the KK mode $l$ decays with gravitational strength to lighter elements of the tower and so the partial decay width is shown to be
$\Gamma^{l}_{l^{\prime}l^{\prime\prime}}\sim m_{l}^{3}/M_{p}^{2}\,.$ (14)
In principle, there are $(m_{l}/m_{\rm KK})^{2}$ possibilities for the $l^{\prime}l^{\prime\prime}$ pair of KK modes that $l$ can decay to, but because of a small violation of KK quantum number, the choice effectively becomes $m_{l}\delta/m_{\rm KK}$. Taking into account the phase factor which is roughly the velocity of decay products at threshold,
$v\sim\sqrt{m_{\rm KK}\ \delta/m_{l}}\,,$ (15)
the total decay width of graviton $l$ is found to be,
$\Gamma^{l}_{\rm tot}$ $\sim$ $\sum_{l^{\prime}<l}\ \ \sum_{0<l^{\prime\prime}<l-l^{\prime}}\Gamma^{l}_{{l^{\prime}}{l^{\prime\prime}}}$ (16)
$\sim$ $\beta^{2}\frac{m_{l}^{3}}{M_{p}^{2}}\times\frac{m_{l}}{m_{\rm KK}}\ \delta\times\sqrt{\frac{m_{\rm KK}\ \delta}{m_{l}}}$
$\sim$ $\beta^{2}\ \delta^{3/2}\frac{m_{l}^{7/2}}{M_{p}^{2}m_{\rm KK}^{1/2}}\,,$
where $\beta$ is parameter that controls the strength of the intra-tower decay amplitudes which correlates with the amplitudes on inhomogeneities in the dark dimension Gonzalo:2022jac . Assuming that for times larger than $1/\Gamma^{{}_{l}}_{\rm tot}$ dark matter which is heavier than the corresponding $m_{l}$ has already decayed, it is straightforward to see that
$m_{l}\sim\left(\frac{M_{p}^{4}\ m_{\rm KK}}{\beta^{4}\ \delta^{3}}\right)^{1/7}t^{-2/7}\,,$ (17)
where $t$ indicates the time elapsed since the big bang.
Now, KK decays impart momentum to the daughter particles, injecting kinetic energy into virialized self-gravitating dark-matter halos. These momentum “kicks” – which have been shown in (15) to depend on $\delta$ — gradually reduce the central density of cuspy dark matter halos and deplete the amount of mass in the halos Peter:2010jy ; DES:2022doi . This process also makes sub-halos that have a virial velocity smaller than the velocity kick more susceptible to tidal disruption as they orbit within a host halo, leading to the suppression of the abundance of low-mass halos and sub-halos at late times. To preserve the structure and stability of galaxy halos, the analysis of Law-Smith:2023czn ; Obied:2023clp severely constrains the $(\tilde{\lambda},\delta,\beta)$ parameter space. The relatively narrow total decay width obtained from substituting the marginalized upper limit on $\delta$ into (16) is substantially smaller than the total decay width obtained in (13). As a consequence, the limit from neutron start heating given in (12) cannot be avoided, and so a dark matter explanation in terms of KK gravitons grows into tension with (6). Only if the KK tower collapses at once, (6) can be made consistent with observations.
Two possible scenarios emerge:
  • The graviton tower decays into the radion zero mode, which could be ultralight and thereby serve as a fuzzy dark matter candidate Anchordoqui:2023tln .
  • On the other side of the coin dark matter is not necessarily related to the KK tower, which still decays to the lightest mode “fast enough” for (6) to be consistent with experiment.
In the Sec. V we give an example of the latter.
IV Constraints on Neutrino Towers
The smallness of the neutrino mass may be ascribed to the fact that right-handed ($R$) neutrinos could live in the bulk Dienes:1998sb ; Arkani-Hamed:1998wuz ; Dvali:1999cn . The coupling of $R$-neutrinos to the left-handed SM neutrinos living on the brane is inversely proportional to the square-root of the bulk volume. The Dirac neutrino mass can be shown to be
$m_{\nu}\sim y\ \langle H\rangle\ \Lambda_{\rm sp}/M_{p}\,,$ (18)
where $y$ is the Yukawa coupling and $\langle H\rangle=246~{}{\rm GeV}$ is the Higgs vacuum expectation value. Note that for $\Lambda_{\rm sp}/M_{p}\ll 1$, the neutrino mass scale is well below the electroweak symmetry breaking scale $\langle H\rangle$, even for order-one Yukawa couplings.
If $R$-neutrinos live in the bulk their KK modes would mix with SM neutrinos modifying their oscillation pattern. The phenomenology is similar to that of light sterile neutrinos, which induce observable short-baseline neutrino oscillations and small perturbations to the neutrino oscillations in solar, atmospheric, and long-baseline experiments that are well-described by standard three-neutrino mixing. The non-observation of these effects in neutrino-detection facilities set 90% C.L. bounds on the size of the largest extra dimension
$R_{\rm KK}<\left\{\begin{array}[]{ll}0.2~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm NO}\\ 0.1~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm IO}\end{array}\right.\,,$ (19)
for normal (NO) and inverted (IO) ordering of neutrino masses Davoudiasl:2002fq ; Machado:2011jt ; Forero:2022skg . However, these bounds could be significantly relaxed in the presence of bulk neutrino masses, because the mixing of the first KK modes to active SM neutrinos becomes suppressed, making the contribution of heavier KK modes to oscillations relatively more important Carena:2017qhd ; Anchordoqui:2023wkm .
The cosmological evolution of $R$-neutrino towers was discussed elsewhere Anchordoqui:2024xvl , with the inference that for a large violation of KK momentum conservation, the $n=2$ scenario with dark dimensions of micron scale remains consistent with observations.
V Six-Dimensional Black Holes
as All Dark Matter
It is well-known that black holes are expected to have formed when very large density perturbations collapse Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx . An attractive idea is that these overdensities are of inflationary origin Hodges:1990bf ; Carr:1993aq ; Ivanov:1994pa . Once the overdense regions come back into causal contact after inflation, they instantly collapse if they exceed a medium-specific threshold. These PBHs fulfill all of the necessary requirements to be a captivating dark matter candidate: they are cold, non-baryonic, quasi-stable, and can be formed in the right abundance to be the dark matter. Actually, if PBHs contribute to more than 10% of the dark matter density, then their energy density today is of the same order as that of the baryons Wu:2021gtd . This cosmic coincidence might hint at a mutual origin for PBHs and the baryon asymmetry of the Universe. Strikingly, PBHs fostered by a transition from a slow-roll to ultraslow-roll during single field inflation can contribute as a significant dark matter component Leach:2001zf , while such a transition of the inflationary background can trigger successful baryogenesis via the Afleck-Dine mechanism Affleck:1984fy .
PBHs evaporate by emitting Hawking radiation, with a temperature inversely proportional to its mass Hawking:1974rv ; Hawking:1975vcx . Nevertheless, a PBH could be stable on cosmological scales and its lifetime could even be longer than the age of the Universe. For example, if the initial mass of a 4D PBH is $M_{\rm BH}\gtrsim 10^{15}~{}{\rm g}$ it can survive until today Page:1976df . Even so, an all-dark-matter interpretation in terms of PBHs have been ruled out through observations across most mass ranges Carr:2020xqk ; Green:2020jor ; Villanueva-Domingo:2021spv . Constraints mostly come from microlensing surveys and non-observation of Hawking radiation. PBHs in the mass range $10^{17.5}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$ remain a viable explanation for all dark matter.
Higher-dimensional black holes of Schwarzschild radii smaller than a micron are: bigger, colder, and longer-lived than a usual 4D black hole of the same mass Argyres:1998qn . In a series of recent publications we have used these black hole properties to extend the PBH range as the entirety of dark matter compared to that in the 4D theory by several orders of magnitude in the low mass window Anchordoqui:2022txe ; Anchordoqui:2022tgp ; Anchordoqui:2024akj ; Anchordoqui:2024dxu . In what follows we particularize our investigation to 6D black holes.
Shortly after being formed a higher-dimensional PBH would emit SM particles and gravitons on the brane as well as gravitons into the bulk. However, the recoil effect due to graviton emission imparts the black hole a relative kick velocity with respect to the brane, allowing the PBH to wander off into the bulk Frolov:2002as ; Frolov:2002gf ; Flachi:2005hi ; Flachi:2006hw ; Anchordoqui:2024jkn . The Hawking evaporation time characterizing the lifetime of such a bulk PBH is estimated to be
$\tau_{\rm BH}\sim r_{s}\ S_{\rm BH}\sim\frac{1}{\Lambda_{\rm sp}}\ \left(\frac{M_{\rm BH}}{\Lambda_{\rm sp}}\right)^{(n+3)/(n+1)}\,,$ (20)
where
$r_{s}\sim\frac{1}{\Lambda_{\rm sp}}\ \left(\frac{M_{\rm BH}}{\Lambda_{\rm sp}}\right)^{1/(n+1)}$ (21)
is the $(4+n)$-dimensional Schwarzschild radius Tangherlini:1963bw ; Myers:1986un and
$S_{\rm BH}\sim(M_{\rm BH}/\Lambda_{\rm sp})^{(n+2)/(n+1)}$ (22)
is the black hole entropy Anchordoqui:2001cg . All in all, taking $n=2$ and $\Lambda_{\rm sp}\sim 10~{}{\rm TeV}$ we obtain
$\tau_{\rm BH}\sim 13.7\left(\frac{M_{\rm BH}}{10^{7.8}~{}{\rm g}}\right)^{5/3}~{}{\rm Gyr}\,.$ (23)
We conclude that the mass range for 6D PBHs to make all the cosmological dark matter is then
$10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}\,.$ (24)
In closing, we stress that if there were 6D primordial near-extremal black holes in nature, then it would be possible to lower the minimum mass allowing a PBH all-dark-matter interpretation, because near-extremal black holes are colder and longer-lived Anchordoqui:2024akj . Note, however, that near-extremal black holes cannot escape from the brane into the bulk if they carry e.g., electromagnetic charge.
VI Ultra-High-Energy Cosmic Rays
       as Probes of Two Dark Dimensions
Extensive air showers induced from ultra-high-energy cosmic rays (UHECRs) provide a window into understanding the most energetic interactions Anchordoqui:2018qom . To be specific, the highest energy cosmic rays interact in the Earth’s atmosphere through collisions with center-of-mass energies $\sqrt{s}\sim 300~{}{\rm TeV}$, viz. not only well above possibilities at the Large Hadron Collider, but even beyond the projections for the FCC-hh.
The initial interaction of a cosmic ray in the atmosphere produces a set of secondary particles (mostly $\pi$’s) carrying a fraction of the primary energy. The $\pi^{0}$’s produced in the first interaction ($\sim 1/3$ of all pions in accord with isospin invariance) promptly decay to a pair of gamma rays. The gamma rays produce $e^{\pm}$ pairs when passing near nuclei. The electrons and positrons re-generate gamma rays via bremsstrahlung, thereby building an electromagnetic cascade. The $\pi^{\pm}$, which carry $2/3$ of the energy lost by the cosmic ray, begin to move through the atmosphere and either decay or interact generating new sets of secondaries. In these interactions the $\pi^{\pm}$ again carry $2/3$ of the parent energy and $1/3$ goes into electromagnetic cascade. After $j$ hadronic cascade generations, only $(2/3)^{j}$ of the total energy remains in the hadronic cascade. The $\pi^{\pm}$ produce air shower muons when they decay. The number of shower muons depends on the amount of energy that is left in the hadronic cascade when pion energies have dropped to the level where decay is more likely than collision. If this happens after relatively few cascade generations, then copious muon production occurs. If the reduction of pion energies takes relatively many generations, then more of the energy will have been lost from the hadronic cascade to the electromagnetic cascade, and meager muon production occurs. Eventually, the electromagnetic cascade dissipates about 90% of the primary cosmic ray energy and the remaining 10% is carried by muons and neutrinos. The total number of muons produced by a cosmic ray nucleus scales roughly as
$N_{\mu}\propto A\ (E/A)^{0.93}\,,$ (25)
where $A$ is the nucleus baryon number Anchordoqui:2018qom . The expectation is then that a cosmic ray nucleus to produce about $A^{0.07}$ more muons than a proton. This means that a shower initiated by an iron nucleus produces about 30% more muons than a proton shower.
Cosmic ray observations suggest that the number of registered muons at the surface of the Earth is by some tens of percentage points higher than expected in simulations. In particular, Auger data suggest that the hadronic component of showers (initiated by cosmic rays of energy $10^{9.8}<E/{\rm GeV}<10^{10.2}$) contains about 30% to 60% more muons than expected PierreAuger:2016nfk . Actually, the muon deficit in air shower simulations with respect to data has been observed by several experiments at different energies Albrecht:2021cxw . The significance of the discrepancy between data and model prediction has been experimentally established at $8\sigma$. The onset of the discrepancy is seen in showers with $E\sim 10^{7.7}~{}{\rm GeV}$, which corresponds to a first interaction at $\sqrt{s}\sim 10~{}{\rm TeV}$ in the nucleon-nucleon system. The relative muon deficit increases approximately linearly with $\ln E$ above the shower energy threshold. The fact that the number of shower stages above this energy also increases linearly with $\ln E$ points towards a collective effect.
A thought-provoking observation is that for two dark dimensions of micron size, we would expect a qualitative change in the evolution of UHECR showers with primary interactions at $\sqrt{s}\gtrsim 10~{}{\rm TeV}$. Scattering processes above the species scale necessarily have internal structure in the compact space and to a certain extent would lead to the production of KK modes. [^1] Moreover, black holes are expected to be produced in collisions with $\sqrt{s}\gg\Lambda_{\rm sp}$ Banks:1999gd . One caveat is that QCD cross sections dominate over the black hole production cross section by a factor of roughly $10^{9}$ Anchordoqui:2003jr . This implies that black holes produced in cosmic ray air showers are effectively unobservable. This is also the case for the production of string massive modes Cornet:2001gy . Pivotally for our purposes, the phase space seen by the emitted graviton grows quite quickly with energy on very general grounds. Namely, for production of KK gravitons carrying energy $\varepsilon$, the scattering cross section is proportional to the multiplicity factor $g_{n}(\varepsilon)$ counting the number of KK gravitons contributing to the process. The mass of each KK mode is the modulus of its momentum in the direction transverse to the brane
$m_{l}^{2}=(\vec{l}/R_{\rm KK})^{2}\,,$ (26)
with $\vec{l}=(l_{1},l_{2},\cdots,l_{n})$. At fixed $\varepsilon$, all gravitons with $m_{l}\leq\varepsilon$ contribute. The multiplicity factor is then estimated to be the volume of an $n$-dimensional sphere with radius $l_{\rm max}=R_{\rm KK}\ \varepsilon$, and is given by
$g_{n}(\varepsilon)=\frac{2\ \pi^{n/2}}{n\Gamma(n/2)}\left(R_{\rm KK}\ \varepsilon\right)^{n}\,,$ (27)
where $\Gamma(x)$ is the Gamma function. As expected, (27) satisfies $g_{0}=1$.
Figure 1: Feynman diagrams for a bremsstrahlung emission of KK gravitons by electrons in the static electric field generated by nuclei.
Next, we propose a new model to address the muon puzzle. To convey our idea, we consider the bremsstrahlung emission of KK gravitons by electrons in the static electric field generated by atmospheric nuclei. The cross section for the diagrams shown in Fig. 1 rises fast, as a power in energy
$\sigma\propto\left(\sqrt{s}/\Lambda_{\rm sp}\right)^{2+n}\,,$ (28)
and while it is unimportant at energies below the species scale, asymptotically it would become the dominant process; e.g., for $n=2$, we have $\sigma\propto s^{2}$ Kazanas:2001ep . Unmodified, this behavior by itself eventually violates unitarity. Eikonal unitarization schemes introduce considerable uncertainty in the eventual energy behavior of the cross section Anchordoqui:2000uh . It is, however, tempting to speculate that at extremely high center-of-mass energies, characterizing the scattering of the first generations of particles in UHECR showers, the cross section for KK production can compete with electromagnetic process, but it is well below QCD cross sections driving the hadronic collisions. If this were the case, then current simulations of extensive air showers would underestimate the ratio of the visible energy dissipated by the hadronic cascade $E_{\rm had}$ to that in the electromagnetic cascade $E_{\rm EM}$. [^2] A point worth noting at this juncture is that photon-KK and photon-axion mixing could produce a similar effect Deffayet:2000pr . Obviously, an underestimation of the ratio $E_{\rm had}/E_{\rm EM}$ would alter the energy reconstruction of the shower. To accomodate the imbalance between $E_{\rm had}$ and $E_{\rm EM}$, the cosmic ray energy $E$ must be slightly increased, and via (25) this implies an increase in the number of muons that could account for the discrepancy between simulations and experiment. [^3] For a vertical transversal of the atmosphere, $\pi^{\pm}$’s below the critical energy $\sim 115~{}{\rm GeV}$ are more likely to decay than to interact again Gondolo:1995fq . Note that the energy of the resulting secondary $\mu^{\pm}$ would be too low to excite the KK tower.
While it is difficult to be precise, two micron-size dark dimensions would predict deviations in the evolution of cosmic ray showers at an energy scale where discrepancies between model building and experimental data have been observed. Ultimately, the impact of new scattering process on cosmic ray air showers, including those that spread out into a large internal space, can be best studied using a full-scale high-fidelity simulation.
VII Conclusions
The motivation for our work is the observation that two extra dimensions on the micron scale can simultaneously address the cosmological and gauge hierarchy problems. To be precise, the dark dimension scenario is not solving the problem, why there is a small cosmological constant, but it connects it to other hierarchies, here the gauge hierarchy problem, which is related to TeV scale physics. Consistency with astrophysical and cosmological observations requires two extra dimensions which do not admit isometries, whereby conservation of the extra dimensional momentum is violated to a large extent, allowing the massive KK modes of the graviton to promptly decay to other lighter graviton modes. Obviously, this prevents the KK towers of gravitons and neutrinos to be dark matter. Instead, we have shown that either a fuzzy radion or else PBHs with masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$ could make all cosmological dark matter.
Above and beyond searches of KK signals at future particle colliders, a variety of astrophysical observations will be able to probe the ideas discussed in this paper. On the one hand, the coldest neutron star (PSR J2144-3933) observed with the Hubble Space Telescope has an upper limit on its effective temperature of about $3.3\times 10^{4}~{}{\rm K}$, but this effective temperature could be constrained down to $2\times 10^{4}~{}{\rm K}$ with JWST-NIRCam, $1.5\times 10^{4}~{}{\rm K}$ at ELT-MICADO, and $9\times 10^{3}~{}{\rm K}$ with TMT-IRIS Raj:2024kjq . The increase sensitivity could naturally have an impact on the upper limit given in (12), providing a test for the KK dynamical dark matter model. On the other hand, the asteroid-mass window for PBHs to be all dark matter will be tested in the near future by microlensing of X-ray pulsars Bai:2018bej ; Tamta:2024pow . This window include the interesting range ($10^{16.5}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{17.5}$) where an all dark matter interpretation in terms of 4D Schwarzschild PBHs is excluded by the non-observation of their Hawking radiation Anchordoqui:2024tdj .
In addition, the long-baseline Deep Underground Neutrino Experiment (DUNE) will probe relevant aspects of neutrino mixing in scenarios with large extra dimensions Berryman:2016szd ; Siyeon:2024pte . Finally, after the due commissioning period, AugerPrime full efficiency data taking (a.k.a. Auger Phase II) started in 2024 and is foreseen to add ten more years of data PierreAuger:2016qzd . These data might help elucidate whether the class of models introduced herein, featuring new-physics processes leading to missing energy in the electromagnetic shower (e.g. via production of KK modes), are at the center of the cosmic-ray muon puzzle.
Acknowledgements
We like to thank Ivano Basile and Gia Dvali for useful discussions. The work of L.A.A. is supported by the U.S. National Science Foundation (NSF Grant PHY-2412679). I.A. is supported by the Second Century Fund (C2F), Chulalongkorn University. The work of D.L. is supported by the Origins Excellence Cluster and by the German-Israel-Project (DIP) on Holography and the Swampland. L.A.A. thanks the Theoretical Physics Laboratory at the UNLP in Argentina for hospitality during completion of this work.